Which of the following is a coterminal angle to the angle 37 degrees?
398
757
143
-350

Answer by greenestamps(13215)



Any angle with a measure that is a multiple of 360 degrees greater than or less than the given angle.

Any positive multiple of 360 added to 37 degrees will be positive and will have degree measure with units digit 7; any negative multiple of 360 added to 37 degrees will be negative and will have a degree measure with a units digit of 3.

Without doing any actual arithmetic, 757 is the only number in the list of choices that satisfies either of those conditions.

ANSWER: 757 (degrees)

CHECK: 37+2(360)=37+720=757